Solve using elimination.
3x + 5y = -8
-2x + 2y = 16

Answer by ikleyn(52802)

.
Multiply first equation by 2 (both sides).

Multiply the second equation by 3.   You will get


 6x + 10y = -16
-6x +  6y =  48


Add the equations.  You will get


      16y = 32


Hence, y = 32/16 = 2.


Substitute the found value y= 2 into the second equation of the original system. You will get


-2x + 2*2 = 16,

-2x = 16 - 2*2 = 12

  x = 12/(-2) = -6.


ANSWER.  x= -6,  y= 2.

Solved.

You may check the answer by substituting the found values into the original equations.